# Retrieving the hot CGM physics from the X-ray radial profile from eROSITA with an IlustrisTNG-based forward model

### v1

@@ Abstract
Aims. Recent eROSITA measurements of the radial profiles of the hot CGM in the Milky-Way stellar mass (MW-mass) regime provide us with a new benchmark to constrain the hot gas around MW-mass central and satellite galaxies and their halo mass distributions. Modelling this rich data set with state-of-the-art hydrodynamical simulations is required to further our understanding of the shortcomings in the current paradigm of galaxy formation and evolution models.
Methods. We develop forward models for the stacked X-ray radial surface brightness profile measured by eROSITA around MW-mass galaxies. Our model contains two emitting components: hot gas (around central galaxies and around satellite galaxies hosted by more massive halos) and X-ray point sources (X-ray binaries and Active Galactic Nuclei). We model the hot gas profile using the TNG300-based products. We generate mock observations with our TNG300-based model (matching stellar mass and redshift with observations) with different underlying halo mass distributions. Therefore, we test the CGM properties as a function of their host halo mass distribution. The point sources are described by a simple point-spread-function (PSF) of eROSITA, and we fit their normalization in this work. In total, we fit the X-ray surface brightness profile with two free parameters: the normalization of satellites in more massive host halos and the normalization of the mean point source emission.
Results. We show that for the same mean stellar mass, a factor 2$\sim$ 2$\times$ 2 × increase in the mean value of the underlying halo mass distribution results in 4$\sim$ 4$\times$ 4 × increase in the stacked X-ray luminosity from the hot CGM. Using empirical models to derive a permissible range of AGN and XRB luminosities in the MW-mass X-ray galaxy stack, we choose our forward model best describing the hot CGM for the eROSITA observations. Our chosen model in the MW stellar mass bin is in good agreement with previous literature results. We find that at $\lesssim 40$ kpc from the galaxy centre, the hot CGM from central galaxies and the X-ray point sources emission (from XRB and AGN) each account for $40-50\%$ of the total X-ray emission budget. Beyond $\sim 40$ kpc, we find that the hot CGM around satellites (probing their more massive host halos with mean $M_{\rm 200m}\sim 10^{14}\ {\rm M}_{\odot}$) dominate the stacked X-ray surface brightness profile.
Conclusions. The gas physics driving the shape of the observed hot CGM (in stellar-mass-selected X-ray stacking experiments) is tightly correlated by the underlying halo-mass distribution. This work provides a novel technique to constrain the AGN X-ray luminosity jointly with the radial hot CGM gas distribution within the halo using measurements from X-ray galaxy stacking experiments. Implementing this technique on other state-of-the-art simulations will provide a new ground for testing different galaxy formation models with observations.
@@ 1 Introduction
The circumgalactic medium (CGM) plays a crucial role in a galaxy’s evolution by directly tracing inflows and outflows of gas driven by various gravitational and non-gravitational mechanisms; see Faucher-Giguère & Oh (2023) for an overview. The non-gravitational mechanisms, such as stellar and Active Galactic Nuclei (AGN) feedback, heat and cool the gas, impacting the evolution of star formation in the galaxy (Donahue & Voit 2022). The relative contributions from shock heating of the gas due to gravitational infall, stellar and AGN heating is model dependent in the current paradigm of galaxy formation and evolution. Particularly, the impact of stellar and AGN feedback depends on the host halo mass, where the former affects the halo masses $\lesssim 10^{12}\ {\rm M}_{\odot}$, and the latter affects the $\gtrsim 10^{12}\ {\rm M}_{\odot}$, respectively (Wechsler & Tinker 2018). The pivotal point, where the relative contributions from stellar and AGN feedback are equally important, occurs at halo mass scales similar to our Milky Way (MW), defining it as a crucial testing range for the models.
@@
The interplay between feedback mechanisms and the galactic atmosphere results in a multiphase CGM, with the hot phase ($T\gtrsim 10^{6}$ K) being typically the most massive and volume-filling component (see review by Tumlinson et al. 2017). The hot CGM radiates in the soft X-ray within the $0.2-2$ keV energy band for MW-mass halos due to thermal hot gas emission. There are various techniques to probe the hot CGM in X-rays via absorption (Galeazzi et al. 2007; Bhattacharyya et al. 2023; Mathur et al. 2023; Wijers et al. 2020; Bogdán et al. 2023) and emission (Koutroumpa et al. 2007; Bertone et al. 2010; van de Voort 2013; Bogdán et al. 2013a, b; Anderson et al. 2016; Bogdán et al. 2017; Li et al. 2017; Das et al. 2019; Zhang et al. 2022; Ponti et al. 2023; Locatelli et al. 2024; Zheng et al. 2024). In particular, emission studies, combined with stacking, allow us to map the large-scale extent of the hot CGM to the halo’s viral radius (Anderson et al. 2015; Oppenheimer et al. 2020; Comparat et al. 2022; Chadayammuri et al. 2022; Zhang et al. 2024a).
@@
Given the advent of eROSITA (Merloni et al. 2024), there have been several studies exploiting the unprecedented statistics for stacking the X-ray emission at the position of optically selected galaxies, such as Comparat et al. (2022); Chadayammuri et al. (2022) and most recently, Zhang et al. (2024a, hereafter Z24). Z24 represent the state-of-the-art hot CGM measurements for MW-mass galaxies given the largest optical galaxy sample with the German half-sky eROSITA coverage in X-rays. They stack $415,627$ galaxies with photometric redshifts, Fullphot, from the DESI Legacy Survey DR 9 (Dey et al. 2019; Zou et al. 2019, 2022) and $30,825$ central galaxies with spectroscopic redshifts from the SDSS DR7 Main Galaxy Sample (Strauss et al. 2002; Abazajian et al. 2009). The latter, the SDSS-based central galaxy sample, is advantaged with spectroscopic information, allowing for classifying galaxies into centrals and satellites with halo mass information (Tinker 2021). Therefore, they retrieve the X-ray surface brightness profile from the hot CGM by empirically modelling the impact due to satellites and AGN and XRB emissions. However, the former, the DESI Legacy survey-based galaxy sample, can not be classified into centrals and satellites, given the limitations in photometric redshift, making the modelling of this dataset challenging. To exploit the highest signal-to-noise data (factor of $13.5$ times more statistics than the spectroscopic sample) to date, in this work, we embark on constructing a forward model to disentangle the hot CGM radial profiles from the X-ray stack of optically selected galaxies.
@@
Among the dominating sources of contamination in X-ray stacking experiments at MW-mass galaxies are (1) the AGN and X-ray binaries (XRB) population of galaxies (Biffi et al. 2018; Vladutescu-Zopp et al. 2023), and (2) the effect of having satellite galaxies in the stacking sample, where these satellites contribute to the averaged X-ray stack with their more massive host galaxy within which they are embedded in the Large Scale Structure (see, e.g., Shreeram et al. 2024; Weng et al. 2024). Given that we are modelling a photometric galaxy sample in this work, we do not have the classification of galaxies into centrals and satellites due to limitations in the (photometric) redshift accuracy for the galaxies in large optical surveys. Therefore, we use our TNG-based forward model fit for the magnitude of the satellite galaxy contribution by setting their normalization of the surface brightness profile free when fitting to data. The contaminating effect of the satellites was quantified in detail in Shreeram et al. (2024), where we find that this effect becomes increasingly significant in the stacking sample as the stellar mass decreases. When conducting blind X-ray stacking analysis at the positions of optically selected galaxies, where centrals and satellites are unclassified, the inclusion of satellites implies that the total measured X-ray surface brightness profile comprises (i) the intrinsic hot gas emission around truly central galaxies and (ii) the contamination of hot gas emission measured around satellites. We emphasize that the latter does not correspond to the emission intrinsic to the satellites as the more massive host (central) galaxy in the vicinity of the satellite dominates the emission, resulting in a negligible contribution of the intrinsic satellite emission (see, e.g., Rohr et al. 2024).
@@
This paper presents a forward model for the stacked galaxy profile comprising the X-ray emitting gas and point source emission. We use the lightcone built with TNG300 in Shreeram et al. (2024) to construct mock galaxy catalogues representing the observations. From our TNG-based mock galaxy catalogue, we predict the hot gas CGM profile contribution to the X-ray galaxy stack from central and satellite galaxies. We parameterize the normalization of the hot gas contribution from satellites, $\mathcal{N}_{\rm sat}$. The hot gas prediction for central galaxies from our forward model with TNG is left unchanged. As for the point-like source contributions from AGN and XRB, we do not use the simulations, which are quite uncertain in their predicted instantaneous AGN activity, but rather we describe their contribution with a normalization parameter $\mathcal{N}_{\rm ps}$ and fix the radial shape using the PSF of eROSITA. This leaves us with two fitting parameters in our model: the contribution of point sources in the stack $\mathcal{N}_{\rm ps}$ and the normalization of the satellite X-ray surface brightness profile $\mathcal{N}_{\rm sat}$. We fit our forward model to the mean X-ray surface brightness profile reported Z24, obtained by stacking MW-mass galaxies. We obtain the point source luminosity obtained from our fitting analysis, and we interpret our results by comparing them with independent empirical models of AGN luminosity functions.
@@
The paper is organized as follows. Sec. 2 expands on the observational data used for modelling in this study. Sec. 3 describes the forward model built in this work; details on the point source and hot gas component are provided in Sec. 3.1 and Sec. 3.2, respectively, and generation of mock galaxy catalogues in Sec. 3.3. Sec. 4 interprets and discusses the results, and Sec. 5 summarizes the main findings of this work.
@@ 2 Data
This study uses the results from Z24, where they use optical data from the Legacy Survey Data release 9 (Dey et al. 2019), and X-ray data from the first four SRG/eROSITA All-Sky Surveys (eRASS:4) within the western galactic hemisphere (Merloni et al. 2024). Here, we summarize the relevant details from Z24 to motivate the forward-model setup, as described in Sec. 3.
The LS DR9 overlaps by $9,340$ deg2 with the western Galactic hemisphere. Z24 define the Fullphot sample containing $1,677,909$ galaxies, which is based on the LS DR9 galaxy catalog from Zou et al. (2019, 2022), where they provide the galaxy properties. The stellar masses, $M_{\star}$, with uncertainties of $\sim 0.2$ dex are provided in the range $11.59.5<\log M_{\star}<$ 11.5 and the photometric redshifts, $z_{\rm phot}$, with uncertainty $\Delta z_{\rm phot}\lesssim 0.03$, spans the range $0.40.01<z_{\rm phot}<$ 0.4 (see Tab. 3 in Z24 for further classification into stellar mass bins). We focus here on the modelling of the X-ray emission from MW-like galaxies, defined by the stellar mass bin $11.010.5<\log_{10}M_{\star}<$ 11.0, and located in the redshift range $0.170.02<z_{\rm phot}<$ 0.17; containing $415,627$ galaxies and median stellar mass and redshift of $5.5\times 10^{10}\ {\rm M}_{\odot}$ and $z_{\rm phot}=0.14$, respectively.
The details on the data analysis pipeline to obtain the observed X-ray surface brightness profiles, based on the LS DR9 (Fullphot) galaxies, are provided in Z24. They stack the X-ray data in different stellar mass bins following the method from Comparat et al. (2022). The first step is to generate X-ray event cubes around every galaxy in the sample within $3$ Mpc. The events within the cube are assigned a physical radial distance from the source, $R_{\rm kpc}$, along with the exposure time, $t_{\rm exp}$, effective collecting area, $A_{\rm eff}$, and the rest frame energy of the event around the source $E_{\rm rest}$. The X-ray surface brightness profile around the galaxy is obtained, with additional correction factors due to absorption and area loss if masking of sources is applied (see Eq. 1 in Z24). We use the background subtracted profiles for modelling in this work (for details on the background treatment in the data, see Appendix A in Z24).
@@ 3 Forward model built in this work
where ${\mathbf{S}}_{\rm X,\ hot\ gas}[r]$ is the X-ray contribution to $\mathbf{S}_{\rm X,\ total}(r)$ from hot gas (further detailed in Sect. 3.2) and ${\mathbf{S}}_{\rm X,\ point-source}[r]$ is the X-ray contribution from AGN and XRB (Sect. 3.1). We then introduce the forward models constructed in this work in Sec. 3.3.
@@ 3.1 The point source component
${\mathbf{S}}_{\rm X,\ point-source}(r)=\mathcal{N}_{\rm ps}\ {\rm PSF}(z,r)$ (2)
@@
where $\mathcal{N}_{\rm ps}$ is the normalization of the point-source component and ${\rm PSF}[z,r]$ is the shape of the point-source component as defined by the survey-averaged eROSITA PSF. Here, the shape of the mean PSF is obtained by converting from angular to physical scale [kpc] using the redshifts of the galaxies in the stacking sample. We constrain $\mathcal{N}_{\rm ps}$ by fitting the observations.
@@
We do not use the TNG300 outputs to predict the contributions from AGN and XRB for the following reasons. As presented in Habouzit et al. (2019), the X-ray Luminosity Function (XLF) for AGN in TNG shows an overproduction of faint AGN at $z=0$, a common problem in hydrodynamical cosmological simulations due to poorly resolved sub-grid feedback prescriptions (Sijacki et al. 2015; Volonteri et al. 2016; Rosas-Guevara et al. 2016; Biffi et al. 2018). Additionally, in TNG, the bright end underpredicts the XLF compared with observations due to the over-efficient kinetic mode of feedback prescriptions with TNG (Habouzit et al. 2019). Given these discrepancies between the hydrodynamical simulations and observations, using the TNG-based predictions for the X-ray AGN contribution would be unreliable for the purpose of this work. As for XRB, whose prediction depends on the star formation rates (SFR) of galaxies (more details in Sec. 4.1), using the TNG SFR values for the mock galaxies to represent observations requires extreme care, given how sensitive the SFR is to the definition of quenched galaxies, the physical apertures used for measurement of the SFR, and the mass resolution of the simulation (Donnari et al. 2019). Additionally, the quenched fraction of satellite galaxies is overestimated in the stellar mass range of $M_{\star}\in 10^{10-11}\ {\rm M}_{\odot}$ at $z=0$ (Donnari et al. 2021), which entails the mass range of interest in this work. These caveats impede us from reliably predicting meaningful estimates for point source contributions intrinsic to the TNG model that can be directly compared with observations.
@@
In this work, we use independent empirical estimates (Sec. 4.1 and Sec. 4.2 ) to predict the allowed range of the mean X-ray luminosity from XRB and AGN, respectively. Since the estimates from these empirical models are independent of TNG, we use them to inform our forward models for the permitted values of point source luminosities, as shown in Fig. 2.
@@
Figure 1: Forward models constructed in this work for the hot CGM from central galaxies by varying the underlying halo mass distribution. Left panel: The purple halo mass distribution (Model 1) is obtained from the mock central galaxy catalogue - constructed with the TNG300 lightcone (LC-TNG300) from Shreeram et al. (2024) - for the X-ray stack from Zhang et al. (2024a) that uses optically detected galaxies with photometric redshifts (Fullphot) from LS DR9 (Dey et al. 2019). Note that the mock galaxy catalogue is generated by matching LC-TNG300 to Fullphot in stellar mass and redshift (see details in Sec. 3.3); the median stellar mass and redshift of the Fullphot (and our mock catalogues) are $5.5\times 10^{10}\ {\rm M}_{\odot}$ and $0.14$, respectively. The underlying halo mass distribution of the Fullphot optical dataset is unknown. The pink distribution (Model 2; with mean $M_{\rm 200m}=5.4\times 10^{12}\ {\rm M}_{\odot}$) discards the top $10\%$ most massive halos before the generation of the mock galaxy catalogue. The yellow distribution (Model 3; with mean $M_{\rm 200m}=3.5\times 10^{12}\ {\rm M}_{\odot}$) discards the top $30\%$ most massive halos before the generation of the mock galaxy catalogue. Right panel: The corresponding X-ray surface brightness profiles in the $0.5-2$ kev energy band (for details on their generation see Sec. 3.2) for the three mock galaxy catalogues with different halo mass distributions, which are shown in the left panel. The profiles are convolved with the eROSITA PSF and they represents the Fullphot dataset in the stellar mass and redshift plane. Nevertheless, due to the impact of the underlying halo mass distribution, the shape and normalization of the hot CGM profiles are impacted, where discarding the most massive halos from the underlying halo distribution results in steeper profiles with lower normalizations.

### v2

# Retrieving the hot circumgalactic medium physics from the X-ray radial profile from eROSITA with an IlustrisTNG-based forward model

@@ Abstract
Aims. Recent eROSITA measurements of the radial profiles of the hot circumgalactic medium (CGM) in the Milky Way stellar mass (MW-mass) regime provide us with a new benchmark to constrain the hot gas around MW-mass central and satellite galaxies and their halo mass distributions. Modeling this rich data set with state-of-the-art hydrodynamical simulations is required to further our understanding of the shortcomings in the current paradigm of galaxy formation and evolution models.
Methods. We developed forward models for the stacked X-ray radial surface brightness profile measured by eROSITA around MW-mass galaxies. Our model contains two emitting components: hot gas (around central galaxies and around satellite galaxies hosted by more massive halos) and X-ray point sources (X-ray binaries (XRBs) and active galactic nuclei (AGNs)). We modeled the hot gas profile using the TNG300-based products. We generated mock observations with our TNG300-based model (matching stellar mass and redshift with observations) with different underlying halo mass distributions. Therefore, we tested the CGM properties as a function of their host halo mass distribution. The point sources are described by a simple point spread function of eROSITA, and we fit their normalization in this work. In total, we fit the X-ray surface brightness profile with two free parameters: the normalization of satellites in more massive host halos and the normalization of the mean point source emission.
Results. We show that for the same mean stellar mass, a factor $\sim 2\times$ increase in the mean value of the underlying halo mass distribution results in $\sim 4\times$ increase in the stacked X-ray luminosity from the hot CGM. Using empirical models to derive a permissible range of AGN and XRB luminosities in the MW-mass X-ray galaxy stack, we choose our forward model that best describes the hot CGM for the eROSITA observations. Our chosen model in the MW stellar mass bin is in good agreement with previous literature results. We find that at $\lesssim 40$ kpc from the galaxy center, the hot CGM from central galaxies and the X-ray point sources emission (from XRBs and AGNs) each account for $40-50\%$ of the total X-ray emission budget. Beyond $\sim 40$ kpc, we find that the hot CGM around satellites (probing their more massive host halos with mean $M_{\rm 200m}\sim 10^{14}\ {\rm M}_{\odot}$) dominates the stacked X-ray surface brightness profile.
Conclusions. The gas physics driving the shape of the observed hot CGM (in stellar-mass-selected X-ray stacking experiments) is tightly correlated with the underlying halo-mass distribution. This work provides a novel technique to constrain the AGN X-ray luminosity jointly with the radial hot CGM gas distribution within the halo using measurements from X-ray galaxy stacking experiments. Implementing this technique on other state-of-the-art simulations will provide a new ground for testing different galaxy formation models with observations.
@@ 1 Introduction
The circumgalactic medium (CGM) plays a crucial role in a galaxy’s evolution by directly tracing inflows and outflows of gas driven by various gravitational and nongravitational mechanisms (see Faucher-Giguère and Oh (2023) for an overview). The nongravitational mechanisms, such as stellar and active galactic nucleus (AGN) feedback, heat and cool the gas, impacting the evolution of star formation in the galaxy (Donahue and Voit 2022). The relative contributions from shock heating of the gas due to gravitational infall, stellar, and AGN heating are model-dependent in the current paradigm of galaxy formation and evolution. Particularly, the impact of stellar and AGN feedback depends on the host halo mass, where the former affects the halo masses $\lesssim 10^{12}\ {\rm M}_{\odot}$, and the latter affects the $\gtrsim 10^{12}\ {\rm M}_{\odot}$, respectively (Wechsler and Tinker 2018). The pivotal point, where the relative contributions from stellar and AGN feedback are equally important, occurs at halo mass scales similar to the one of our Milky Way (MW), defining it as a crucial testing range for the models.
@@
The interplay between feedback mechanisms and the galactic atmosphere results in a multiphase CGM, with the hot phase ($T\gtrsim 10^{6}$ K) typically being the most massive and volume-filling component (see review by Tumlinson et al. 2017). The hot CGM radiates in the soft X-ray within the 0.220.2-2 keV energy band for MW-mass halos due to thermal hot gas emission. There are various techniques to probe the hot CGM in X-rays via absorption (Galeazzi et al. 2007; Bhattacharyya et al. 2023; Mathur et al. 2023; Wijers et al. 2020; Bogdán et al. 2023) and emission (Koutroumpa et al. 2007; Bertone et al. 2010; van de Voort 2013; Bogdán et al. 2013a, b; Anderson et al. 2016; Bogdán et al. 2017; Li et al. 2017; Das et al. 2019; Zhang et al. 2022; Ponti et al. 2023; Locatelli et al. 2024; Zheng et al. 2024). In particular, emission studies, combined with stacking, allow us to map the large-scale extent of the hot CGM to the halo’s virial radius (Anderson et al. 2015; Oppenheimer et al. 2020; Comparat et al. 2022; Chadayammuri et al. 2022; Zhang et al. 2024a).
@@
Given the advent of eROSITA (Merloni et al. 2024), there have been several studies exploiting the unprecedented statistics for stacking the X-ray emission at the position of optically selected galaxies, such as Comparat et al. (2022), Chadayammuri et al. (2022), and most recently Zhang et al. (2024a, hereafter Z24). Z24 represent the state-of-the-art hot CGM measurements for MW-mass galaxies, given the largest optical galaxy sample with the German half-sky eROSITA coverage in X-rays. They stack $415,627$ galaxies with photometric redshifts, Fullphot, from the DESI Legacy Survey DR 9 (Dey et al. 2019; Zou et al. 2019, 2022) and $30,825$ central galaxies with spectroscopic redshifts from the SDSS DR7 Main Galaxy Sample (Strauss et al. 2002; Abazajian et al. 2009). The latter, the SDSS-based central galaxy sample, has the advantage of spectroscopic information, allowing for galaxies to be classified into centrals and satellites with halo mass information (Tinker 2021). Therefore, Z24 retrieve the X-ray surface brightness profile from the hot CGM by empirically modeling the impact due to satellites, AGNs, and XRB emissions. However, the former, the DESI Legacy survey-based galaxy sample, cannot be classified into centrals and satellites, given the limitations in photometric redshift, making the modeling of this dataset challenging. To exploit the highest signal-to-noise data (a factor of $13.5$ times more statistics than the spectroscopic sample) to date, in this work, we embark on constructing a forward model to disentangle the hot CGM radial profiles from the X-ray stack of optically selected galaxies.
@@
Among the dominating sources of contamination in X-ray stacking experiments at MW-mass galaxies are (1) the AGN and X-ray binary (XRB) population of galaxies (Biffi et al. 2018a; Vladutescu-Zopp et al. 2023), and (2) the effect of having satellite galaxies in the stacking sample, where these satellites contribute to the averaged X-ray stack with their more massive host galaxy within which they are embedded in the large-scale structure (see e.g., Shreeram et al. 2025; Weng et al. 2024). Given that we are modeling a photometric galaxy sample in this work, we do not have the classification of galaxies into centrals and satellites due to limitations in the (photometric) redshift accuracy for the galaxies in large optical surveys. Therefore, we use our TNG-based forward model fit for the magnitude of the satellite galaxy contribution by setting the normalization of the surface brightness profile free when fitting to data. The contaminating effect of the satellites was quantified in detail in Shreeram et al. (2025), where we find that this effect becomes increasingly significant in the stacking sample as the stellar mass decreases. When conducting blind X-ray stacking analysis at the positions of optically selected galaxies, where centrals and satellites are unclassified, the inclusion of satellites implies that the total measured X-ray surface brightness profile comprises (i) the intrinsic hot gas emission around truly central galaxies and (ii) the contamination of hot gas emission measured around satellites. We emphasize that the latter does not correspond to the emission intrinsic to the satellites as the more massive host (central) galaxy in the vicinity of the satellite dominates the emission, resulting in a negligible contribution of the intrinsic satellite emission (see e.g., Rohr et al. 2024).
@@
This paper presents a forward model for the stacked galaxy profile comprising the X-ray emitting gas and point source emission. We use the lightcone built with TNG300 in Shreeram et al. (2025) to construct mock galaxy catalogs representing the observations. From our TNG-based mock galaxy catalog, we predict the hot gas CGM profile contribution to the X-ray galaxy stack from central and satellite galaxies. We parameterize the normalization of the hot gas contribution from satellites, $\mathcal{N}_{\rm sat}$. The hot gas prediction for central galaxies from our forward model with TNG is left unchanged. As for the point-like source contributions from AGNs and XRBs, we do not use the simulations, which are quite uncertain in their predicted instantaneous AGN activity, but instead we describe their contribution with a normalization parameter, $\mathcal{N}_{\rm ps}$, and fix the radial shape using the point spread function (PSF) of eROSITA. This leaves us with two fitting parameters in our model: the contribution of point sources in the stack, $\mathcal{N}_{\rm ps}$, and the normalization of the satellite X-ray surface brightness profile, $\mathcal{N}_{\rm sat}$. We fit our forward model to the mean X-ray surface brightness profile reported Z24, obtained by stacking MW-mass galaxies. We obtain the point source luminosity from our fitting analysis, and we interpret our results by comparing them with independent empirical models of AGN luminosity functions.
@@
The paper is organized as follows. Sec. 2 expands on the observational data used for modeling in this study. Sec. 3 describes the forward model built in this work; details on the point source and hot gas component are provided in Sec. 3.1 and Sec. 3.2, respectively, and details on the generation of mock galaxy catalogs in Sec. 3.3. Sec. 4 interprets and discusses the results, and Sec. 5 summarizes the main findings of this work.
@@ 2 Data
This study uses the results from Z24, where they use optical data from the Legacy Survey Data release 9 (Dey et al. 2019), and X-ray data from the first four SRG/eROSITA All-Sky Surveys (eRASS:4) within the western galactic hemisphere (Merloni et al. 2024). Here, we summarize the relevant details from Z24 to motivate the forward-model setup, as is described in Sec. 3.
The LS DR9 overlaps by $9,340$ deg2 with the western Galactic hemisphere. Z24 define the Fullphot sample containing $1,677,909$ galaxies, which is based on the LS DR9 galaxy catalog from Zou et al. (2019, 2022), where they provide the galaxy properties. The stellar masses, $M_{\star}$, with uncertainties of $\sim 0.2$ dex are provided in the range $9.5<\log M_{\star}<11.5$ and the photometric redshifts, $z_{\rm phot}$, with uncertainty $\Delta z_{\rm phot}\lesssim 0.03$, spans the range $0.01<z_{\rm phot}<0.4$ (see Tab. 3 in Z24 for further classification into stellar mass bins). We focus here on the modeling of the X-ray emission from MW-like galaxies, defined by the stellar mass bin $10.5<\log_{10}M_{\star}<11.0$, and located in the redshift range $0.02<z_{\rm phot}<0.17$; containing $415,627$ galaxies and median stellar mass and redshift of $5.5\times 10^{10}\ {\rm M}_{\odot}$ and $z_{\rm phot}=0.14$, respectively.
The details on the data analysis pipeline to obtain the observed X-ray surface brightness profiles, based on the LS DR9 (Fullphot) galaxies, are provided in Z24. They stack the X-ray data in different stellar mass bins following the method from Comparat et al. (2022). The first step is to generate X-ray event cubes around every galaxy in the sample within $3$ Mpc. The events within the cube are assigned a physical radial distance from the source, $R_{\rm kpc}$, along with the exposure time, $t_{\rm exp}$, effective collecting area, $A_{\rm eff}$, and the rest frame energy of the event around the source, $E_{\rm rest}$. The X-ray surface brightness profile around the galaxy is obtained, with additional correction factors due to absorption and area loss if masking of sources is applied (see Eq. 1 in Z24). We use the background-subtracted profiles for modeling in this work (for details on the background treatment in the data, see Appendix A in Z24).
@@ 3 Forward model built in this work
where ${\mathbf{S}}_{\rm X,\ hot\ gas}[r]$ is the X-ray contribution to $\mathbf{S}_{\rm X,\ total}(r)$ from hot gas (further detailed in Sect. 3.2) and ${\mathbf{S}}_{\rm X,\ point-source}[r]$ is the X-ray contribution from AGNs and XRBs (Sect. 3.1). We then introduce the forward models constructed in this work in Sec. 3.3.
@@ 3.1 The point source component
${\mathbf{S}}_{\rm X,\ point-source}(r)=\mathcal{N}_{\rm ps}\ {\rm PSF}(z,r),$ (2)
@@
where $\mathcal{N}_{\rm ps}$ is the normalization of the point-source component and ${\rm PSF}[z,r]$ is the shape of the point-source component as defined by the survey-averaged eROSITA PSF. Here, the shape of the mean PSF was obtained by converting from angular to physical scale [kiloparsecs] using the redshifts of the galaxies in the stacking sample. We constrained $\mathcal{N}_{\rm ps}$ by fitting the observations.
@@
We did not use the TNG300 outputs to predict the contributions from AGNs and XRBs for the following reasons. As is presented in Habouzit et al. (2019), the X-ray luminosity function (XLF) for AGNs in TNG shows an overproduction of faint AGNs at $z=0$, a common problem in hydrodynamical cosmological simulations due to poorly resolved sub-grid feedback prescriptions (Sijacki et al. 2015; Volonteri et al. 2016; Rosas-Guevara et al. 2016; Biffi et al. 2018a). Additionally, in TNG, the bright end underpredicts the XLF compared with observations due to the over-efficient kinetic mode of feedback prescriptions with TNG (Habouzit et al. 2019). Given these discrepancies between the hydrodynamical simulations and observations, using the TNG-based predictions for the X-ray AGN contribution would be unreliable for the purpose of this work. As for XRBs, whose prediction depends on the star formation rates (SFRs) of galaxies (more details in Sec. 4.1), using the TNG SFR values for the mock galaxies to represent observations requires extreme care, given how sensitive the SFR is to the definition of quenched galaxies, the physical apertures used for measurement of the SFR, and the mass resolution of the simulation (Donnari et al. 2019). Additionally, the quenched fraction of satellite galaxies is overestimated in the stellar mass range of $M_{\star}\in 10^{10-11}\ {\rm M}_{\odot}$ at $z=0$ (Donnari et al. 2021), which entails the mass range of interest in this work. These caveats impede us from reliably predicting meaningful estimates for point source contributions intrinsic to the TNG model that can be directly compared with observations.
@@
In this work, we use independent empirical estimates (Sec. 4.1 and Sec. 4.2) to predict the allowed range of the mean X-ray luminosity from XRBs and AGNs, respectively. Since the estimates from these empirical models are independent of TNG, we use them to inform our forward models for the permitted values of point source luminosities, as is shown in Fig. 2.
@@
Figure 1: Forward models constructed in this work for the hot CGM from central galaxies by varying the underlying halo mass distribution. Left panel: Purple halo mass distribution (Model 1) obtained from the mock central galaxy catalog – constructed with the TNG300 lightcone (LC-TNG300) from Shreeram et al. (2025) – for the X-ray stack from Zhang et al. (2024a) that uses optically detected galaxies with photometric redshifts (Fullphot) from LS DR9 (Dey et al. 2019). Note that the mock galaxy catalog is generated by matching LC-TNG300 to Fullphot in stellar mass and redshift (see details in Sec. 3.3); the median stellar mass and redshift of the Fullphot (and our mock catalogs) are $5.5\times 10^{10}\ {\rm M}_{\odot}$ and $0.14$, respectively. The underlying halo mass distribution of the Fullphot optical dataset is unknown. The pink distribution (Model 2; with mean $M_{\rm 200m}=5.4\times 10^{12}\ {\rm M}_{\odot}$) discards the top $10\%$ most massive halos before the generation of the mock galaxy catalog. The yellow distribution (Model 3; with mean $M_{\rm 200m}=3.5\times 10^{12}\ {\rm M}_{\odot}$) discards the top $30\%$ most massive halos before the generation of the mock galaxy catalog. Right panel: Corresponding X-ray surface brightness profiles in the 0.520.5-2 kev energy band (for details on their generation see Sec. 3.2) for the three mock galaxy catalogs with different halo mass distributions, which are shown in the left panel. The profiles are convolved with the eROSITA PSF, and they represent the Fullphot dataset in the stellar mass and redshift plane. Nevertheless, due to the impact of the underlying halo mass distribution, the shape and normalization of the hot CGM profiles are impacted, where discarding the most massive halos from the underlying halo distribution results in steeper profiles with lower normalizations.